OpenStudy (anonymous):

How do you write a polynomial function, with the numbers, -1,0 and -4?

OpenStudy (anonymous):

not sure what you mean - are these numbers the zeros of the polynomial/

OpenStudy (anonymous):

Sorry here are the full instructions, Write a polynomial function of least degree with integral coefficients that has the given zeros.

OpenStudy (anonymous):

oh ok x(x+ 1)(x+4) expand: = x(x^2 + 5x + 4) = x^3 + 5x^2 + 4x

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

yw